# Airborne Sense and Detect of Drones using Deep Learning and LiDAR Point Clouds

### v2

@@ 3 Detection Model
We modify the PointPillars model for air-to-air drone scenarios by extending the scan space vertically. Furthermore, we introduce a novel sparse convolution algorithm to accelerate the detection model’s backbone layer.
3.1 Original PointPillars
Figure 1: The PointPillars architecture consists of three key building blocks: Pillar Feature Net, Backbone, and Detection Head. The lower section visualizes binary images of various stages of the backbone in the latent space.
@@ 3.1 Original PointPillars
PointPillars is a deep learning-based object detection algorithm originally designed for autonomous driving applications on ground. It can operates on 3D point clouds generated by LiDAR sensors. As shown in Fig. 1, the PointPillars network consists of three main components: the voxelization layer (pillar feature net), the backbone layer, and the detection head. The voxelization layer takes raw point cloud data as input and converts it into a sparse 3D voxel grid representation. The voxel grid is organized into pillars, which are vertical stacks of voxels that preserve the spatial and geometric information of the original point cloud data. The feature dimension of each point is denoted by $D$, encompassing nine dimensions: the spatial coordinates $x$, $y$ and $z$; reflectance r; the pillar centre offset values $x_{p}$ and $y_{p}$; and the distance to the arithmetic mean of the pillar $x{{}_{c}}$, $y_{c}$ and $z_{c}$. Points within each pillar are first processed through a 1D convolution block. The maximum values from each pillar are then taken as features and scattered to their corresponding positions on the 2D pseudo-image of the pillar grid. The backbone layer is a convolutional neural network (CNN) that takes as input the features of the points in each pillar and extracts high-level semantic features that capture the spatial and semantic characteristics of the objects in the point cloud data. The backbone layer is the most computationally expensive part of the detection model. The detection head takes as input the features extracted by the backbone layer and generates a set of predicted 3D bounding boxes for each object in the point cloud data. The detection head is a two-stage network that first predicts the objectness score for each voxel in the point cloud data and then refines the location and size of each predicted box.
3.2 Adaptations to Drone Scenarios
@@
In autonomous driving applications the LiDAR sensor is mounted on top of a vehicle and the point cloud data is collected from a ground-level perspective. In airborne sense and detect, the LiDAR sensor observes the full 3D space, with objects of interest frequently appearing above and below the horizontal plane of the drone-mounted LiDAR. Our use case (maintenance inspection survey) involves flight below the 120m ceiling (due to regulations) and is typically less than 60m to meet the data resolution and ground sample density requirements of the survey specification. An example of a LiDAR point cloud (accumulated) generated by an airborne survey with a drone is shown in Fig. 2. While the original network was trained to detect multiple classes of object (cars, pedestrians, etc.) we focused on a single object class and annotated only observations of drones in the point cloud data.
Figure 2: Reconstructed LiDAR point clouds generated by an airborne survey with a drone
@@
In the PointPillars architecture, the 3D space is divided into a set of pillars, and each pillar corresponds to a 2D slice of the 3D space. The anchors are predefined 3D bounding boxes that are used to classify and regress the objects within each pillar. By default, the PointPillars algorithm uses a fixed-size anchor boxes and z-center of anchor boxes is usually fixed at a constant value, since the elevation of the object being detected is assumed to be relatively constant in the local area covered by the LiDAR sensor. PointPillars performs object localization by first regressing the object’s position in the X-Y plane, followed by additional regression steps to determine height and elevation.
@@
To better detect airborne objects such as drones with varying elevation, we made the following adaptation to the original PointPillars. We horizontally stratify the LiDAR field of view into layers where each layer extends to 1m in height (typically 10 layers above and 10 layers below the elevation of the LiDAR). We allocate an anchor box of size 1.6m×1.6m×1.0m (length × width × height) to each layer, with the centre of the anchor box located at the centre of each layer. We also assign different class names to each layer, such as from $drone\_0$, $drone\_1$, …, $drone\_20$. Each class has a corresponding anchor box. The optimal elevation prediction will be related to both the z-centre localization accuracy and the classification accuracy. As shown in the original PointPillars, the localization regression residual along the z-axis is given by Equation 1, where $z^{gt}$ and $z^{a}$ are the z-centre of ground truth box and anchor box respectively, $h^{a}$ is the height of the anchor box. The object classification loss is given by Equation 2, where $p^{a}$ is the class probability of an anchor, $\alpha$ and $\gamma$ are parameters used to adjust the loss function in order to address class imbalance,
@@
here $\alpha$ and $\gamma$ are set to constants of 0.25 and 2 respectively as in the original PointPillars.
This approach allows the algorithm to better detect objects with varying elevation, such as a drone. The size and spacing of the layers are determined based on the characteristics of the drone being detected. We use the same loss functions introduced in the original PointPillars.
@@ 3.3.1 Sparse convolution with scatter operation
We describe the convolution with scatter operation first. With standard convolution, at each output location, the product between each element of the filter and the input element it overlaps with is computed and the results are summed to obtain the output in the current location, see Fig. 3(a) and 3(b). However, in standard convolution, observe that the same input element multiplied by different filter weights contributes to different output elements and the location of those output elements neighbor each other. For example, in Fig.3(b), the input element $x_{13}$ is multiplied by different filter weights and the results are added to different output locations. Specifically, $x_{13}\times w_{9}$ is added to the location of $y_{7}$, $x_{13}\times w_{8}$ is added to the location of $y_{8}$ and so on. We can implement the convolution in an alternative way: multiply each input element with different filter weights and add the result to each corresponding output location which is decided by the filter weight location, see Fig. 3(c). We call this implementation Convolution with Scatter Operation (Unlike standard convolution in which each neighboring pixel is multiplied by the corresponding filter weight and the result is added to the final output position, in this proposed convolution, each pixel is multiplied by different filter weights and the results are added to different output positions. Therefore we call it the Convolution with Scatter Operation).
@@
For the input image with $C$ channel, the convolution with scattering can be expressed as follows:
@@
The psudo code of the convolution with scatter operation is given in Algorithm 1. It is clear the number of multiplications and additions required is the same as standard convolution, it is $O(p\times q\times k^{2})$. For sparse convolution with scatter operation, the time complexity is $O(l\times k^{2})$ where $l$ is the number of active sites and $l\ll p\times q$.
@@
for $i\leftarrow 0$ to $p-1$ do
       for $j\leftarrow 0$ to $q-1$ do
             for $m\leftarrow 0$ to $k-1$ do
                   for $n\leftarrow 0$ to $k-1$ do
@@ 3.3.2 GPU implementation
In our GPU implementation of Sparse Convolution with Scatter Operation, we use matrix multiplication to perform the channel-wise multiplication and summation of the input and a filter kernel. To achieve a higher performance, we first use the GPU prefix sum [42] to compress the active sites in the input image into a dense format. Positions of all active sites are recorded in a 2D mask array, where an element of the 2D array is set to 1 if the corresponding site is an active site, otherwise it is set to 0. The prefix sum will then be applied to this 2D mask array. NVIDIA cuBLAS library [43], a library of highly optimized matrix multiplication routines, is used to perform the channel-wise multiplication and summation. The results are then scattered to the corresponding output positions according to the scattering rule. To avoid write conflicts, we use the atomicAdd() function [44] to perform the final addition. Unlike the method in [27], we do not apply any data lowering in our GPU implementation. Additionally, there is no the rule book or hash table used in our GPU implementation, as rule book or hash table increases the memory overhead and the computation complexity.
In this work, we present the initial implementation of the Sparse Convolution with Scatter Operation. Our GPU implementation follows gather-GEMM-scatter dataflow [29], which needs to read and write DRAM at least three times, respectively. There is significant room for optimization using the following techniques: 1) perform computation on on-chip SRAM (shared memory) using tiling, which reduces access to DRAM memory, and 2) overlap memory access with computation.
@@ 4 Model Training
    The operational site is mapped by a drone mounted LiDAR. The LiDAR is oriented in forward looking mode to capture point cloud data for the monitoring scenario. During the mapping flights, multiple drones are flown in controlled patterns within the segregated airspace and under supervision. The observations of these drones in the LiDAR point cloud data are labelled using the recorded position and navigation data of each drone in the swarm to generate an annotated training dataset. With the help of GPS correction techniques such as Real Time Kinematics (RTK) and Post Processed Kinematics (PPK), we achieve a high-quality annotation in which the error can be controlled $<5cm$.
@@
    The trained model is embedded on the drone and deployed for tactical monitoring of drones within the operating theatre.
@@ 4.1 Training Data Collection and Labelling
Figure 4: (a) Livox Avia non-repetitive scanning pattern (100ms frame) (b) The wireframe mesh model of the M300, which is defined by 207409 triangular facets.
@@
L1 data exchange is supported by converting the raw LiDAR data into the interoperable LAS format (v1.2) and Smoothed Best Estimate (SBET) files for 3D reconstruction in EGM96 geoid. The observations of swarm drones can then be labelled with bounding boxes calculated from the recorded flight logs. For each point, the LAS (Point Record Format 3) records the relative coordinate X, Y, Z and reflectivity along with the GNSS timestamp at microsecond precision. The SBET file provides the timestamped 6 degree-of-freedom LiDAR pose information throughout the flight enabling transformation of the LAS point data between body, earth, and PointPillars coordinate systems. In the case of the real-time stream from the Avia, the data is already in body fixed frame aligned to the PointPillars coordinate system. Example frames (100ms frame) with swarm drone labelled are shown in Fig. 5.
@@ 4.2 Data Augmentation
Our augmentation strategy instead uses a scenario Digital Twin to accurately model the relationship between the relative positions of the LiDAR and observed drones and the heterogeneous spatial and temporal resolution of the LiDAR sensor. The simulated points generated by the Digital Twin exhibit a sampling pattern and sensing distortion that matches its relative location within the LiDAR field of view. Fig. 6 shows the simulated drone cluster generated by ray tracing a triangulated model of the M300 drone at different locations within the LiDAR field of view. Simulation of 100ms frames takes an average 0.2075ms on an 8 GPU server. This involves ray tracing a drone model with 207,409 triangles. In effect the simulator can generate data at 481 times the productivity of the real LiDAR. Analysis of the approach is presented in the results section. Fig. 7 shows some examples of simulated drones which are inserted in real background. The number of points of each drone is larger than 10 and all drone positions are randomly selected.
The Digital Twin enables the development cycle to virtualize the cost and risk of collecting data in safety critical scenarios such as violations of swarm separation rules and incursions by uncooperative drones.
Figure 6: Synthetic clusters of drones generated by ray tracing

### v3

@@ 3 Detection Model
In this section, we propose an airborne PointPillars model that meets the requirements of scan space extension and the real-time processing needs of airborne applications.
3.1 Standard PointPillars
@@
PointPillars is an object detection algorithm, originally applied to automotive hazard perception on the ground. It can operate on 3D point clouds generated by LiDAR sensors. As shown in Fig. 1, the PointPillars network consists of three main components: the voxelization layer (pillar feature net), the backbone layer, and the detection head. The voxelization layer takes raw point cloud data as input and converts it into a sparse 3D voxel grid representation. The voxel grid is organized into pillars, which are vertical stacks of voxels that preserve the spatial and geometric information of the original point cloud data. The feature dimension of each point is denoted by $D$, encompassing nine dimensions: the spatial coordinates $x$, $y$ and $z$; reflectance $r$; the pillar center offset values $x_{p}$ and $y_{p}$; and the distance to the arithmetic mean of the pillar $x{{}_{c}}$, $y_{c}$ and $z_{c}$. Points within each pillar are first processed through a 1D convolution block. The maximum values from each pillar are then taken as features and scattered to their corresponding positions on the 2D pseudo-image of the pillar grid. The backbone layer is a convolutional neural network (CNN) that takes as input the features of the points in each pillar and extracts high-level semantic features that capture the spatial and semantic characteristics of the objects in the point cloud data. The backbone layer is the most computationally expensive part, consuming 81.8% of the total computation (see Section 8.1.2). The detection head takes as input the features extracted by the backbone layer and generates a set of predicted 3D bounding boxes for each object in the point cloud data. The detection head predicts the objectness score for each voxel in the point cloud data and then refines the location and size of each predicted box.
3.2 Airborne PointPillars
@@
In autonomous driving applications the LiDAR sensor is mounted on top of a vehicle and the point cloud data is collected from a ground-level perspective. In airborne sense and detect, the LiDAR sensor observes the full 3D space, with objects of interest frequently appearing above and below the horizontal plane of the drone-mounted LiDAR. Our use case (maintenance inspection survey) involves flight at altitudes of up to 120m (due to regulations) but can also extend below the surrounding ground level, such as when descending into the Harburg Lock (see Section 8.2) to conduct inspections within its structural enclosure. In photogrammetry and high-resolution LiDAR surveys, flights often require drones to operate in close-proximity to structures to achieve the necessary ground sample density (GSD), sometimes within just a few meters of the target surface. This increases the complexity of the object detection challenge, as drones must reliably detect and avoid both static infrastructure and other airborne assets in constrained environments. While the standard PointPillars was trained to detect multiple classes of object (cars, pedestrians, etc.), we focused on a single object class and annotated only observations of drones in the point cloud data.
@@
In the PointPillars architecture, the 3D space is divided into a set of pillars, and each pillar corresponds to a 2D slice of the 3D space. The anchors are predefined 3D bounding boxes that are used to classify and regress the objects within each pillar. By default, the PointPillars algorithm uses fixed-size anchor boxes and the z-center of the anchor boxes is usually fixed at a constant value, since the elevation of the object being detected is assumed to be relatively constant in the local area covered by the LiDAR sensor. PointPillars performs object localization by first regressing the object’s position in the x-y plane, followed by additional regression steps to determine height and elevation. This means that if a dataset includes objects at varying altitudes and more anchor boxes are added with different elevations, the PointPillars model can be trained to predict the 3D locations of objects at varying altitudes.
@@
To better detect airborne objects such as drones with varying elevation, we made the following adaptation to the standard PointPillars. We horizontally stratify the LiDAR field of view into layers where each layer extends to 1m in height (typically 10 layers above and 10 layers below the elevation of the LiDAR). We allocate an anchor box of size 1.6m×1.6m×1.0m (length × width × height) to each layer, with the center of the anchor box located at the center of each layer. We also assign different class names to each layer, from drone_0drone\_0, drone_1drone\_1, …, drone_20drone\_20. Each class has a corresponding anchor box. The optimal elevation prediction will be related to both the z-center localization accuracy and the classification accuracy. As shown in the standard PointPillars, the localization regression residual along the z-axis is given by Equation 1, where $z^{gt}$ and $z^{a}$ are the z-center of ground truth box and anchor box respectively, $h^{a}$ is the height of the anchor box. The object classification loss is given by Equation 2, where $p^{a}$ is the class probability of an anchor, $\alpha$ and $\gamma$ are parameters used to adjust the loss function in order to address class imbalance,
@@
here $\alpha$ and $\gamma$ are set to constants of 0.25 and 2 respectively as in the standard PointPillars.
This approach allows the algorithm to better detect objects with varying elevation, such as a drone. The size and spacing of the layers are determined based on the characteristics of the drone being detected. We use the same loss functions introduced in the standard PointPillars.
@@
To evaluate the performance, the model was tested during supervised flights at Newcastle Aerodrome, Ireland (see Appendix D.1) and Hamburg Port, Germany (see Appendix D.2). The two test sites comprise a mix of built and natural environment and provide distinct background scenery from one another. The Newcastle site is a coastal railway bridge over a tidal inlet and is surrounded by rocks, gravel, and sand dunes. The Hamburg site is the Harburg Lock, which is a navigable waterway system that consists of a concrete chamber with gates at either end, surrounded by buildings and trees. Maintenance inspection surveys were performed at both sites involving multiple drones deployed with camera and LiDAR sensors to collect point and pixel datasets for structural analysis. The flight profiles at both sites were also significantly different, in particular the perspective and relative vantage points from which drones were observed in the LiDAR training data. This is significant as the detection model decorates input points to encode the shape, appearance, and relative position of observed objects, while the grid and pillar layout encodes the spatial relationship between the object and its surrounding environment.
@@ 3.3.1 Sparse convolution with scatter operation
We describe the convolution with scatter operation first. With standard convolution, at each output location, the product between each element of the filter and the input element it overlaps with is computed and the results are summed to obtain the output in the current location, see Fig. 2(a) and 2(b). However, in standard convolution, observe that the same input element multiplied by different filter weights contributes to different output elements and the location of those output elements neighbor each other. For example, in Fig.2(b), the input element $x_{13}$ is multiplied by different filter weights and the results are added to different output locations. Specifically, $x_{13}\times w_{9}$ is added to the location of $y_{7}$, $x_{13}\times w_{8}$ is added to the location of $y_{8}$ and so on. We can implement the convolution in an alternative way: multiply each input element with different filter weights and add the result to each corresponding output location which is decided by the filter weight location, see Fig. 2(c). We call this implementation Convolution with Scatter Operation (Unlike standard convolution in which each neighboring pixel is multiplied by the corresponding filter weight and the result is added to the final output position, in this proposed convolution, each pixel is multiplied by different filter weights and the results are added to different output positions. Therefore we call it the Convolution with Scatter Operation).
@@
For the input image with $C$ channels, the convolution with scattering operation can be expressed as follows:
@@
The pseudocode of the convolution with scatter operation is given in Algorithm 1. It is clear the number of multiplications and additions required is the same as standard convolution, it is $O(p\times q\times k^{2})$. For sparse convolution with scatter operation, the time complexity is $O(l\times k^{2})$ where $l$ is the number of active sites and $l\ll p\times q$.
@@
for $i\leftarrow 0$ to p1p-1 do
 for $j\leftarrow 0$ to q1q-1 do
    for $m\leftarrow 0$ to k1k-1 do
       for $n\leftarrow 0$ to k1k-1 do
@@ 3.3.2 GPU implementation
In our GPU implementation of Sparse Convolution with Scatter Operation, we use matrix multiplication to perform the channel-wise multiplication and summation of the input and a filter kernel. To achieve a higher performance, we first use the GPU prefix sum [52] to compress the active sites in the input image into a dense format. Positions of all active sites are recorded in a 2D mask array, where an element of the 2D array is set to 1 if the corresponding site is an active site, otherwise it is set to 0. The prefix sum will then be applied to this 2D mask array. NVIDIA cuBLAS library [53], a library of highly optimized matrix multiplication routines, is used to perform the channel-wise multiplication and summation. The results are then scattered to the corresponding output positions according to the scattering rule. To avoid write conflicts, we use the atomicAdd() function [54] to perform the final addition. Unlike the method in [36], we do not apply any data lowering in our GPU implementation. Additionally, there is no rule book or hash table used in our GPU implementation, as the use of rule book or hash table increases the memory overhead and the computational complexity.
In this work, we present the initial implementation of the Sparse Convolution with Scatter Operation. Our GPU implementation follows the gather-GEMM-scatter dataflow [38], which needs to read and write DRAM at least three times, respectively. There is significant room for optimization using the following techniques: 1) perform computation on on-chip SRAM (shared memory) using tiling, which reduces access to DRAM memory, and 2) overlap memory access with computation.
@@ 4 Model Training
    The operational site is mapped by a drone mounted LiDAR. The LiDAR is oriented in forward looking mode to capture point cloud data for the monitoring scenario. During the mapping flights, multiple drones are flown in controlled patterns within the segregated airspace and under supervision. The observations of these drones in the LiDAR point cloud data are labelled using the recorded position and navigation data of each drone in the swarm to generate an annotated training dataset. With the help of GNSS correction techniques such as Real Time Kinematics (RTK) and Post Processed Kinematics (PPK), we achieve a high-quality annotation in which the error can be controlled <5cm<5cm.
@@
    The trained model is embedded on the drone and deployed for tactical monitoring of drones within the operating theater.
@@ 4.1 Training Data Collection and Labelling
Figure 3: (a) Livox Avia non-repetitive scanning pattern (100ms frame, 24k points) (b) The wireframe mesh model of the M300, which is defined by 207409 triangular facets.
@@
L1 data exchange is supported by converting the raw LiDAR data into the interoperable LAS format (v1.2) and Smoothed Best Estimate (SBET) files for 3D reconstruction in the EGM96 geoid. The observations of swarm drones can then be labelled with bounding boxes calculated from the recorded flight logs. For each point, the LAS (Point Record Format 3) records the relative coordinate X, Y, Z and reflectivity along with the GNSS timestamp at microsecond precision. Leveraging RTK and PPK techniques, it ensures a positional accuracy within $5cm$. The SBET file provides the timestamped 6 degree-of-freedom LiDAR pose information throughout the flight enabling transformation of the LAS point data between body, earth, and PointPillars coordinate systems. In the case of the real-time stream from the Avia, the data is already in body fixed frame aligned to the PointPillars coordinate system. Example frames (100ms frame) with swarm drone labelled are shown in Fig. 4.
@@ 4.2 Data Augmentation
Our augmentation strategy instead uses a scenario Digital Twin to accurately model the relationship between the relative positions of the LiDAR and observed drones and the heterogeneous spatial and temporal resolution of the LiDAR sensor. We incorporate the NVIDIA OptiX Ray Tracing engine [58] to accelerate the simulation of the LiDAR, for details see Appendix A. The simulated points generated by the Digital Twin exhibit a sampling pattern and sensing distortion that matches its relative location within the LiDAR field of view. Fig. 5 shows the simulated drone cluster generated by ray tracing a triangulated model of the M300 drone at different locations within the LiDAR field of view. Simulation of 100ms frames takes an average 0.2075ms on an 8 GPU server. This involves ray tracing a drone model with 207,409 triangles. In effect the simulator can generate data at 481 times the productivity of the real LiDAR. Analysis of the approach is presented in the results section. Fig. 6 shows some examples of simulated drones which are inserted into a real background. The number of points of each drone is greater than 10 and all drone positions are randomly selected. The Digital Twin enables the development cycle to virtualize the cost and risk of collecting data in safety critical scenarios.
Figure 5: Synthetic clusters of drones generated by ray tracing.